Zeronknight (zeronknight):

Larry passed 5 tests and failed 1. Julie passed 8 tests and failed 2. Donna passed 8 tests and failed 3. * What is the order of these three students' ratios of passed to failed from least to greatest? Julie, Donna, Larry Larry, Donna, Julie Donna, Julie, Larry Debbie, Julie, Luke

Zeronknight (zeronknight):

@Mehek14 Help? o:

Mehek (mehek14):

I guess the first step is setting up the ratios?

Mehek (mehek14):

So you would do \(\dfrac{number~passed}{number~failed}\)

Zeronknight (zeronknight):

Larry - 5/1 Julie - 8/2 Donna -8/3

Mehek (mehek14):

Now list them from least to greatest

Zeronknight (zeronknight):

I have no clue? 8/1 5/1 8/3?

Mehek (mehek14):

Turn them into decimals if it helps

Zeronknight (zeronknight):

Idk how lol

Mehek (mehek14):

5/1 = 5 8/2 = 4 8/3 = 2.666. . .

Mehek (mehek14):

Now you can list them :P

Zeronknight (zeronknight):

Oh. LOL. So just divide by what they failed?

Zeronknight (zeronknight):

5/1 8/2 8/3

Mehek (mehek14):

Yeah

Mehek (mehek14):

Passed divided by failed

Mehek (mehek14):

lol switch them, you ordered from greatest to least

Zeronknight (zeronknight):

Oh LOL. 8/3 8/2 5/1

Mehek (mehek14):

Yes So which option is that?

Zeronknight (zeronknight):

3rd one.

Mehek (mehek14):

Yep Idk where they got Debbie and Luke for D o.o

Zeronknight (zeronknight):

o.o